OpenStudy (luigi0210):

Differentiate:

OpenStudy (luigi0210):

\[\LARGE y=\frac{e^x}{1+x}\]

OpenStudy (anonymous):

What class are you in?

OpenStudy (anonymous):

And for this one just use the quotient rule.

OpenStudy (luigi0210):

So.. \[\LARGE \frac{(1+x)(e^x)-(e^x)}{(1+x)^2}\] \[\LARGE y'=\frac{x}{(1+x)^2}\]

OpenStudy (luigi0210):

And I'm in calc 1.. again T_T

OpenStudy (anonymous):

gg

OpenStudy (anonymous):

And check your numerator.

OpenStudy (anonymous):

And I never learned about e^x in calc 1 .. T_T

OpenStudy (luigi0210):

Yea, I'm missing an e^x.. just noticed >.<

OpenStudy (anonymous):

..Which is weird..because it seems like every other school did... T_T

OpenStudy (anonymous):

lol..... T_T

OpenStudy (anonymous):

Can anyone here help me with advanced calc?

OpenStudy (mathmale):

\[\frac{ d }{ dx }\frac{ e ^{x} }{ 1+x }=\frac{ (1+x)e ^{x}-e ^{x} }{ (1+x)^{2} }=\frac{ e ^{x}+xe ^{x}-e ^{x} }{ (1+x)^{2} }\]

OpenStudy (mathmale):

and that can be reduced to \[\frac{ xe ^{^{x}} }{(1+x)^{2} }\]

OpenStudy (mathmale):

So, Luigi, you're missing an e^x.

OpenStudy (mathmale):

(As you yourself said.)

OpenStudy (mathmale):

Anna: Please post your question in the usual way; if someone can help you, you'll likely hear from that person (or persons).

OpenStudy (anonymous):

xe^x/(1+x)^2 as @mathmale answer